### v11

# Valuation rings as limits of complete intersection rings

Abstract.
We show that a valuation ring containing its residue field of characteristic $p>0$ is a filtered colimit of complete intersection ${\bf F}_{p}$-algebras.
Key words : immediate extensions, pseudo convergent sequences, pseudo limits, ultrapowers, smooth morphisms, Henselian Rings.
2020 Mathematics Subject Classification: Primary 13F30, Secondary 13A18,13L05,13B40.
Introduction
The Zariski Uniformization Theory [32] plays an important role in Desingularization Theory.
Theorem 1.
(Zariski) Let $V$ be a valuation ring containing a field $L\supset{\bf Q}$. Then $V$ is a filtered union of its regular local subrings essentially of finite type over $L$.
This theorem says in fact that $V$ is a filtered union of its smooth $L$-subalgebras of finite type and the following theorem is connected with it.
Theorem 2.
[21, Theorem 21] If $V\subset V^{\prime}$ is an immediate extension of valuation rings containing $\bf Q$ then $V^{\prime}$ is a filtered colimit of smooth $V$-algebras.
An inclusion $V\subset V^{\prime}$ of valuation rings is an immediate extension if it is local as a map of local rings and induces isomorphisms between the value groups and the residue fields of $V$ and $V^{\prime}$. We remind that a filtered colimit is a limit indexed by a small category that is filtered (see [29, 002V] or [29, 04AX]). A filtered union is a filtered direct limit in which all objects are subobjects of the final colimit, so that in particular all the transition arrows are monomorphisms.
The above theorem does not hold when the valuation ring extension is algebraic and contains a field of positive characteristic (see [19, Example 3.13] inspired from [17] and [19, Remark 6.10]), but we got the following result.
Theorem 3.
[25, Theorem 4] Let $V\subset V^{\prime}$ be an immediate extension of valuation rings containing a field and $K\subset K^{\prime}$ its fraction field extension. If $K^{\prime}=K(x)$ for some algebraically independent system of elements $x$ over $K$ then $V^{\prime}$ is a filtered union of smooth $V$-subalgebras of $V^{\prime}$.
As in [23, Proposition 18](see also [26, Proposition 1]) we have the following corollary.
Corollary 4.
[25, Corollary 24] Let $V\subset V^{\prime}$ be an immediate extension of valuation rings containing a field and $K\subset K^{\prime}$ its fraction field extension. If $K^{\prime}=K(x)$ for some algebraically independent system of elements $x$ over $K$ and $V$ is Henselian then every finite system of polynomials over $V$, which has a solution in $V^{\prime}$ has also one in $V$.
The ideas of this corollary were used to show some conjectures of M. Artin in [18, Theorems 1.3, 1.4] (see also [10, Theorem 5.3.1]) using the so called the General Néron desingularization (see [18, Theorem 2.5], [28, Theorem 1.1] and [10, Theorem 5.2.56]).
A consequence of Theorem 3 is a form of Zariski’s Uniformization Theorem in a special case of positive characteristic.
Corollary 5.
([25, Corollary 5], [27, Theorem 3]) Let $V$ be a valuation ring containing its residue field $k$ with a value group $\Gamma$ free as a $\bf Z$-module, $\mathrm{val}$ its valuation and $K$ its fraction field. Assume that $K=k(x,y)$ for some algebraically independent elements $x,y$, $x=(x_{i})_{i\in I},y=(y_{j})_{j\in J}$ over $k$ such that $\mathrm{val}(y)$ is a basis in $\Gamma$. Then $V$ is a filtered union of its smooth $k$-subalgebras.
This Corollary says that $V$ is a filtered union of its regular local subrings essentially of finite type over $k$.
In general, we got the following theorem based on Theorem 3 and [24, Theorem 1].
Theorem 6.
([25, Theorem 6], [27, Theorem 17]) Let $V\subset V^{\prime}$ be an immediate extension of valuation rings containing a field. Then $V^{\prime}$ is a filtered union of its complete intersection $V$-subalgebras of finite type.
A complete intersection $V$-algebra of finite type is a $V$-algebra of type $C/(P)$, where $C$ is a polynomial $V$-algebra and $P$ is a regular system of elements of $C$. Thus the above theorem says that $V^{\prime}$ is a filtered union of its $V$-subalgebras of type $C/(P)$. Since $V^{\prime}$ is local it is enough to say that $V^{\prime}$ is a filtered union of its $V$-subalgebras of type $_{h}/(P), TT$ being a polynomial $V$-algebra, $0\not=h\in T$ and $P$ is a regular system of elements of $T$. Clearly, $T_{h}$ is a smooth $V$-algebra and in fact it is enough to say that $V^{\prime}$ is a filtered union of its $V$-subalgebras of type $G/(P)$, where $G$ is a smooth $V$-algebra and $P$ is a regular system of elements of $G$. Conversely, a $V$-algebra of such type $G/(P)$ has the form $T_{h}/(P)$ for some $T,h,P$ using [28, Theorem 2.5]. By abuse we understand by a complete intersection $V$-algebra a $V$-algebra of such type $G/(P)$, or $T_{h}/(P)$ which are not assumed to be flat over $V$.
Like Corollary 5 we have the following corollary.
Corollary 7.
([25, Corollary 7], [27, Corollary 18], see also [26, Theorem 19]) Let $V$ be a valuation ring containing its residue field $k$ with a free value group (for example when it is finitely generated) $\Gamma$. Then $V$ is a filtered union of its complete intersection $k$-subalgebras of finite type.
The goal of this paper is to extend somehow Corollary 5 (see Theorems 14, 15, 19) and to establish the following theorem, that is an extension of the above corollary for the case when $\Gamma$ is not finitely generated.
Theorem 8.
Let $V$ be a valuation ring containing a perfect field $F$ of positive characteristic, $k$ its residue field and $\Gamma$ its value group. Then $V$ is a filtered colimit of complete intersection $F$-algebras of finite type if one of the following conditions holds:
  1. (1)
    $k\subset V$,
  2. (2)
    $\dim V=1$,
  3. (3)
    $V$ is Henselian.
The proof is given in Theorems 24, 28. It is possible that the above theorem could give a complete intersection version of the Desingularization Theory.
When $\Gamma$ is not finitely generated we had to use the cross-sections (these are retractions $\Gamma\rightarrow K^{*}$ of the valuation) as in [31, Lemma 7.9], or [2, 3.3.39, 3.3.40] obtained using some ultrapower methods as we did it in [21] when $V\supset{\bf Q}$. The same methods are necessary to reduce to the case when the residue field $k\subset V$ in the case when $V$ is Henselian (see Theorem 26).
The above theorem connects the theory of valuation rings with the theory of complete intersection local rings. We remind that for a given complete local ring $(R,{\mathfrak{m}},k)$ a minimal free resolution over $R$ of the residue field $k$ is described in [30, Theorem 4] and others as [4, Theorem 2.7], [5, §6], [8, Proposition 1.5.4], [9, §3], [6, Theorem 2.5], [16, Theorem 4.1].
1. Filtered colimits of smooth algebras
A possible extension of Zariski’s Uniformization Theorem is the following theorem.
Theorem 9.
(B. Antieau, R. Datta [1, Theorem 4.1.1]) Every perfect valuation ring of characteristic $p>0$ is a filtered union of its smooth ${\bf F}_{p}$-subalgebras.
Using the above theorem we get a consequence of Theorem 3 similar to Corollaries 5, 7.
Corollary 10.
Let $V\subset V^{\prime}$ be an immediate extension of valuation rings of positive characteristic $p$ and $K\subset K^{\prime}$ their fraction field extension. Assume that $K=K^{p}$ and $K^{\prime}$ is a pure transcendental extension of $K$. Then $V^{\prime}$ is a filtered union of its smooth ${\bf F}_{p}$-subalgebras. If $K^{\prime}/K$ is not a pure transcendental extension then $V^{\prime}$ is a filtered union of its complete intersection ${\bf F}_{p}$-subalgebras of finite type.
For the proof of the first statement, note that $V$ is a filtered union of its smooth ${\bf F}_{p}$-subalgebras by Theorem 9 and $V^{\prime}$ is a filtered union of its smooth $V$-subalgebras by Theorem 3. The proof of the second statement goes similarly using Theorem 6 instead Theorem 3. It is also necessary to note that a complete intersection algebra over a complete intersection algebra is still a complete intersection algebra by Lemma [24, Lemma 6].
We remind below [21, Lemma 7], which is an extension of [12, Proposition 3], and [20, Proposition 5].
Lemma 11.
For a commutative diagram of ring morphisms
[…]
with $B$ finitely presented over $A$, a $b\in B$ that is standard over $A$ (this means a special element from the ideal $H_{B/A}$ defining the non smooth locus of $B$ over $A$, for details see for example [21, Lemma 4]), and a nonzerodivisor $a\in A^{\prime}$ that maps to a nonzerodivisor in $V$ that lies in every maximal ideal of $V$, there is a smooth $A^{\prime}$-algebra $S$ such that the original diagram factors as follows:
[…]
Let $V$ be a valuation ring, $\lambda$ a fixed limit ordinal and $v=\{v_{i}\}_{i<\lambda}$ a sequence of elements in $V$ indexed by the ordinals $i$ less than $\lambda$. Then $v$ is pseudo convergent if $\mathrm{val}(v_{i}-v_{i^{\prime\prime}})<\mathrm{val}(v_{i^{\prime}}-v_{i^{\prime\prime}})\ \ \mbox{for}\ \ i<i^{\prime}<i^{\prime\prime}<\lambda$ (see [11]). A pseudo limit of $v$ is an element $v\in V$ with $\mathrm{val}(v-v_{i})<\mathrm{val}(v-v_{i^{\prime}})\ \ \mbox{(that is,}\ \ \mathrm{val}(v-v_{i})=\mathrm{val}(v_{i}-v_{i^{\prime}}))\ \ \mbox{for}\ \ i<i^{\prime}<\lambda$.
When for any $\gamma\in\Gamma$ it holds $\mathrm{val}(v_{i}-v_{i^{\prime}})>\gamma$ for $i<i^{\prime}<\lambda$ large we call $v$ fundamental. $V$ is complete if every fundamental sequence over $V$ has a pseudo limit in $V$ (in this case it is unique and called limit). Using the fundamental sequences of a valuation ring $V$ we may construct its completion. An immediate extension $V\subset V^{\prime}$ of valuation rings is dense if any element of $V^{\prime}$ is a limit of a fundamental sequence of $V$. Thus if $V\subset V^{\prime}$ is dense then $V,V^{\prime}$ have the same completion.
Next we remind an application of Lemma 11.
Proposition 12.
([21, Proposition 9] For a ring $A$, a dense extension of valuation rings (see Section 3) $V\subset V^{\prime}$, $K$ the fraction field of $V$, a ring morphism $A\rightarrow V$, a finitely presented $A$-algebra $B$, and maps
$A\rightarrow B\rightarrow V\ \ \mbox{such that}\ \ B\rightarrow K\ \ \mbox{factors through some $A$-smooth localization of $B$}$
suppose that there exist a smooth $A$-algebra $S^{\prime}$ and a factorization $A\rightarrow B\rightarrow S^{\prime}\rightarrow V^{\prime}$. Then there exist a smooth $A$-algebra $S$ and a factorization $A\rightarrow B\rightarrow S\rightarrow V$. In particular, there exist a smooth $A$-algebra $S$ and a factorization $A\rightarrow B\rightarrow S\rightarrow V$ if there exist a smooth $A$-algebra $\hat{S}$ and a factorization $A\rightarrow B\rightarrow{\hat{S}}\rightarrow{\hat{V}}$, $^{\hat{V}}$ being the completion of $V$.
Remark 13.
In particular of the above proposition, let $V\subset V^{\prime}$ be a separable dense extension of valuation rings, $K\subset K^{\prime}$ its fraction field extension, $B$ a finitely presented $V$-algebra and $\rho:B\rightarrow V$ a $V$-morphism. As $K^{\prime}/K$ is separable there exists a smooth $V$-algebra $S^{\prime\prime}$ such that $B\xrightarrow{\rho}V\rightarrow K^{\prime}$ factors through $S^{\prime\prime}$. Thus $\rho(H_{B/V})V^{\prime}\not=0$ and so $\rho(H_{B/V})\not=0$ and $V/\rho(H_{B/V})V\cong V^{\prime}/\rho(H_{B/V})V^{\prime}$, since $V\subset V^{\prime}$ is dense. If $B\xrightarrow{\rho}V\rightarrow V^{\prime}$ factors through a smooth $V$-algebra $S^{\prime}$ then $\rho$ factors through a smooth $V$-algebra $S$ using Lemma 11.
The following theorem follows from the above remark, Theorem 3 and the proof of Corollary 5.
Theorem 14.
Let $V\subset V^{\prime}$ be an immediate extension of valuation rings of positive characteristic, $\mathrm{val}$ the valuation of $V^{\prime}$ and $K\subset K^{\prime}$ their fraction field extension. Assume that
  1. (1)
    $V$ contains its residue field $k$ and $x$ is a system of elements of $V$ such that $\mathrm{val}(x)$ forms a $\bf Z$-linear basis of the value group of $V$.
  2. (2)
    $V$ is a separable dense extension of $V\cap k(x)$.
  3. (3)
    $K^{\prime}$ is a separable dense extension of a pure transcendental field extension of $K$.
Then $V^{\prime}$ is a filtered colimit of smooth $k$-algebras.
Proof.
As in the proof of Corollary 5, $V\cap k(x)$ is a filtered union of some localizations of polynomial algebras over $k$. Let $y$ be a transcendental basis of $K^{\prime}$ over $K$ such that $K(y)\subset K^{\prime}$ is dense. The separable dense extensions $V\cap k(x)\subset V$ and $K(y)\subset K^{\prime}$ are filtered colimits of smooth $V\cap k(x)$-algebras, respectively $V^{\prime}\cap K(y)$-algebras by the above remark. It remains to apply Theorem 3 to the extension $V\subset V^{\prime}\cap K(y)$.
Using Theorem 9 we get a similar theorem.
Theorem 15.
Let $V\subset V^{\prime}$ be an immediate extension of valuation rings of positive characteristic $p$, $\mathrm{val}$ the valuation of $V^{\prime}$ and $K\subset K^{\prime}$ their fraction field extension. Assume that
  1. (1)
    There exists a valuation subring $W$ of $V$ such that $W=W^{p}$.
  2. (2)
    $V$ is a separable dense extension of $W$.
  3. (3)
    $K^{\prime}$ is a separable dense extension of a pure transcendental field extension of $K$.
Then $V^{\prime}$ is a filtered colimit of smooth ${\bf F}_{p}$-algebras.
In the case of one dimensional valuation extension there exists a stronger form of Theorem 14, which we will present next together with some necessary facts concerning the so called the possible defectless of a valuation ring.
Let $V$ be a valuation ring with value group $\Gamma$, $K$ its fraction field and as usual its valuation $\mathrm{val}$. Let $K^{\prime}$ be a finite field extension of $K$, $v_{i}$, $1\leq i\leq r$ the valuations of $K^{\prime}$ extending $\mathrm{val}$ and $V_{i}$ be the valuation rings of $K^{\prime}$ defined by $v_{i}$. Let $e_{i},f_{i}$ be the ramification index respectively the degree of the residue field extension of $V_{i}$ over $V$. It is well known that
$K^{\prime}:K]\geq\sum_{i=1}^{r}e_{i}f_{i}.$
If this inequality is equality for all finite field extensions $K^{\prime}$ of $K$ we say that $V$ is defectless (see [14] for details and examples).
Very useful is the following particular form of the Generalized Stability Theorem of F. V. Kuhlmann (see [13, Theorem 1.1], [14, Theorem 5.1]).
Theorem 16.
(Kuhlmann) Let $V\subset V^{\prime}$ be an extension of valuation rings with the same residue field, $\Gamma\subset\Gamma^{\prime}$ its value group extension and $K\subset K^{\prime}$ its fraction field extension. Assume $K^{\prime}/K$ is a finite type field extension and $\Gamma^{\prime}/\Gamma$ is a finitely generated free $\bf Z$-module of rank $\operatorname{tr\,deg}K^{\prime}/K$. If $V$ is defectless then $V^{\prime}$ is defectless too.
Corollary 17.
Let $V$ be a valuation ring of positive characteristic with finitely generated value group $\Gamma$, $\mathrm{val}$ its valuation and fraction field $K$. Assume $V$ contains its residue field $k$ and $K=k(y)$ for some elements $y=(y_{1},\ldots,y_{r})$ of $V$ such that $\mathrm{val}(y_{i})$, $1\leq i\leq r$ is a basis of the free $\bf Z$-module $\Gamma$. Then $V$ is defectless.
For the proof apply the above theorem to $k\subset V$ since $k$ is defectless for the trivial valuation.
The following lemma is [22, Lemma 8] but we give in sketch a proof here.
Lemma 18.
Let $V\subset V^{\prime}$ be an immediate extension of one dimensional valuation rings and $K\subset K^{\prime}$ their fraction field extension. If $V$ is defectless and $K^{\prime}/K$ is algebraic then $K^{\prime}/K$ is dense.
Proof.
We may consider only the case when $K^{\prime}/K$ is finite. Let $K^{h}$, $K^{\prime h}$ be the Henselizations of $K$ respectively $K^{\prime}$ and $L=K^{h}(K^{\prime})\subset K^{\prime h}$. Note that $K^{h}$ is defectless because $K$ is so (see [14, Theorem 2.3]). Then $L:K^{h}]=e_{L/K^{h}}f_{L/K^{h}}=1$ because there exists an unique extension of the valuation of $K^{h}$ to $L$ and so $K^{\prime}\subset L\subset K^{h}$. But $K\subset K^{h}$ is dense because $\dim V=1$.
Theorem 19.
Let $V\subset V^{\prime}$ be an immediate extension of one dimensional valuation rings of positive characteristic, $\mathrm{val}$ the valuation of $V^{\prime}$ and $K\subset K^{\prime}$ their fraction field extension. Assume that
  1. (1)
    $V$ contains its residue field $k$ and $x$ is a finite system of elements of $V$ such that $\mathrm{val}(x)$ forms a $\bf Z$-linear basis of the value group of $V$.
  2. (2)
    There exists an algebraic separable field extension $L\subset K$ of $k(x)$ such that $V$ is a separable dense extension of $V\cap L$.
  3. (3)
    $K^{\prime}$ is a separable dense extension of a pure transcendental field extension of $K$.
Then $V^{\prime}$ is a filtered colimit of smooth $k$-algebras.
Proof.
The proof goes as in Theorem 14, the only thing which we should see is that the extension $L/k(x)$ is dense, which follows from Lemma 18 and Corollary 17.
The following lemma is useful in the next section to reduce the study of filtered colimit of complete intersection algebra to the case of complete valuation rings.
Lemma 20.
Let $V\subset V^{\prime}$ be an extension of valuation rings and $^{\hat{V}^{\prime}}$ the completion of $V^{\prime}$. Assume that the field extension of $V\subset V^{\prime}$ is separable. Then $V^{\prime}$ is a filtered colimit of complete intersection $V$-algebras if $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras. Moreover, if the above fraction field extension is not separable, $E$ is a finitely presented $V$-algebra and $w:E\rightarrow V^{\prime}$ is a $V$-morphism such that $w(H_{E/V})V^{\prime}\not=0$ then $w$ factors through a complete intersection $V$-algebra if $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras.
Proof.
We follow the proofs of [21, Lemma 7, Proposition 9]. Assume that $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras. Let $E$ be a finitely generated $V$-subalgebra of $V^{\prime}$ and $w$ the inclusion $\subset V^{\prime}. EE$ is finitely presented by [15, Theorem 4]. By separability of the fraction field extension of $V\subset E$, the singular locus $w(H_{E/V})V^{\prime}$ has the form $zV^{\prime}$ for some nonzero $z\in V^{\prime}$. The same thing follows when the separability is missing but $w(H_{E/V})V^{\prime}\not=0$ . Actually, changing if necessary $E$ by $E[Z]$ and $w$ by the possible non injective map $E[Z]\rightarrow V^{\prime}$ given by $Z\rightarrow z$ we may assume that $w(H_{E/V})V^{\prime}$ has the form $dV^{\prime}$ for some $d\in E$.
Suppose that $E\cong V[Y]/I$, $Y=(Y_{1}\ldots,Y_{m})$, $I$ being a finitely generated ideal. As in the proof of [21, Proposition 20] we may assume that for some polynomials $f=(f_{1},\ldots,f_{r})$ from $I$, we have $d\in NME$ for some $N\in((f):I)$ and a $r\times r$-minor $M$ of the Jacobian matrix $(\partial f_{i}/\partial Y_{j})$.
Note that $^{\hat{w}}$ the composite map $E\xrightarrow{w}V^{\prime}\subset{\hat{V}}^{\prime}$ factors through a complete intersection $V$-algebra $B$, let us say $\hat{w}$ is the composite map $E\rightarrow B\xrightarrow{\rho^{\prime}}{\hat{V}}^{\prime}$. We may assume that $B$ has the form $T_{h}/(P)$ for a polynomial $V$-algebra $T=V[U]$, $U=(U_{1},\ldots,U_{l})$, $h\in T$ and a regular system of elements $P=(P_{1},\ldots,P_{t})$ from $T$. Suppose the map $\rho^{\prime}$ is given by $U\rightarrow{\hat{u}}$ for some ${\hat{u}}\in({\hat{V}}^{\prime})^{l}$. Thus h(u^)h($\hat${u}}) is invertible. Clearly, the map $\bar{w}$ induced by $w$ modulo $d^{3}$ factors through ${\bar{B}}=B/(d^{3})$, that is $\bar{w}$ is the composite map ${\bar{E}}=E/(d^{3})\rightarrow{\bar{B}}\rightarrow{\bar{V}}^{\prime}=V^{\prime}/(d^{3})\cong{\hat{V}}^{\prime}/(d^{3})$, the last map being given by $U\rightarrow u+d^{3}({\hat{V}}^{\prime})^{l}$, for some $u\in(V^{\prime})^{l}$ which is $\hat{u}$ modulo $d^{3}$. We have $P(u)=d^{3}z$ for some $z\in V^{\prime t}$. Set $D=(V[U,Z]/(P-d^{3}Z))_{h}$, $Z=(Z_{1},\ldots,Z_{t})$ and let $\rho:D\rightarrow V^{\prime}$ be the map $(U,Z)\rightarrow(u,z)$. Note that $D$ is a complete intersection $V$-algebra. As $w$ factors through $\rho$ modulo $d^{3}$ there exists a smooth $D$-algebra $D^{\prime}$ such that $w$ factors through $D^{\prime}$ using Lemma 11 applied for $A=V$ and $A^{\prime}=D$. It follows that $D^{\prime}$ is a complete intersection $V$-algebra as well and it is enough to apply [28, Lemma 1.5].
2. Model Theory of valued fields
Our goal is here to extend Corollary 7.
Corollary 21.
Let $V$ be a valuation ring containing a field $F$ of positive characteristic, $k$ its residue field and $\Gamma$ its value group. Assume that $\Gamma$ is finitely generated, $k/F$ is separably generated (that is there exists a transcendental basis $x$ of $k$ over $F$ such that the field extension $F(x)\subset k$ is algebraic separable) and $V$ is Henselian. Then $V$ is a filtered union of its complete intersection $F$-subalgebras.
Proof.
The conclusion goes from Corollary 7 if we show that $k\subset V$. A lifting of $k$ to $V$ could be done when $V$ is Henselian and char $k=0$ (see [31, Theorem 2.9] but the proof goes in the same way in positive characteristic and $k$ is separably generated over $F$. In particular the lifting could be done when $k/F$ is of finite type.
Corollary 22.
Let $V$ be a valuation ring containing a field $F$ of positive characteristic, $k$ its residue field, $\Gamma$ its value group and $K$ its fraction field. Assume that there exists a cross-section $s:\Gamma\rightarrow K^{*}$ and either $k\subset V$, or $k/F$ is separably generated and $V$ is Henselian. Then $V$ is a filtered union of its complete intersection $F$-subalgebras.
The proof goes as in Corollaries 7, 21 with $V_{0}=V\cap k(s(\Gamma))$. More precisely, $V_{0}$ is a filtered union of $V\cap k(s(\Gamma^{\prime}))$ for all finitely generated subgroups (so free) $\Gamma^{\prime}$ of $\Gamma$, which are localizations of polynomial $k$-algebras.
There are two problems two extend Corollaries 5, 21. These are too reduce to the case:
i) when $V$ has a cross-section $s:\Gamma\rightarrow K^{*}$,
ii) when $V$ contains its residue field.
For i) we may use [31, Proposition 5.4, Lemma 7.9], or [2, 3.3.39, 3.3.40]. In [21, Theorem A.10] there exists the following possible variant of these results.
Theorem 23.
For a valuation ring $V$ with value group $\Gamma$, there is a countable sequence of ultrafilters ${\mathscr{U}}_{1},{\mathscr{U}}_{2},\dots$ on some respective sets $U_{1},U_{2},\dots$ for which the valuation rings $\{V_{n}\}_{n\geq 0}$ defined inductively by $V_{0}:=V$ and $V_{n+1}:=\prod_{{\mathscr{U}}_{n+1}}V_{n}$ are such that the valuation ring
[…]
where $\widetilde{K}$ and $\widetilde{\Gamma}$ are the fraction field and the value group of $\widetilde{V}$.
Theorem 24.
Let $V$ be a valuation ring containing its residue field $k$ of positive characteristic and $\Gamma$ its value group. Then $V$ is a filtered colimit of complete intersection $k$-algebras.
Proof.
Certainly we may assume that $V$ is not a field, otherwise the conclusion is trivial. Let $E$ be a finitely generated $k$-subalgebra (in fact finitely presented by [15, Theorem 4]) of $V$. We will show that $w$ factors through a complete intersection $V$-algebra. It is enough to replace $V$ by the intersection of $V$ with the fraction field $K_{E}$ of $E$.
As $K_{E}$ is a finite type field extension of $F$ we may assume $\Gamma$ is countable generated and $k$ is countable generated over $F$ by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorem 23 for $V$. Note that $\widetilde{k}\subset\widetilde{V}$ because $k\subset V$. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$. Using Theorem 6 we see that $\widetilde{V}$ is a filtered union of its complete intersection $\widetilde{V}_{0}$-subalgebras.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so through a complete intersection $k$-algebra $D$ since $\widetilde{k}/k$ is separable. Thus $w$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$. The proof ends applying [28, Lemma 1.5].
The problem from ii) is also hard. As we explain in the proof of Corollary 21 the lifting of the residue field $k$ of $V$ when the characteristic $p>0$ could be done when $V$ is Henselian and $k/{\bf F}_{p}$ is of finite type. To glue the liftings of finite type subfields of $k$ is another problem which is solved using Model Theory. On the other hand our method uses completions, which are Henselian but in dimension one.
We need [21, Proposition A.6] (useful also in the proof of Theorem 23), which is obtained using [7, Theorem 6.1.4] and says in particular the following:
Proposition 25.
Let $V$ be a valuation ring with value group $\Gamma$ and $\tau=card\ \Gamma$. Then there exists an ultrafilter $\mathcal{U}$ on a set $U$ with $card\ U=\tau$ such that any system of polynomial equations $(g_{i}((X_{j})_{j\in J})_{i\in I}$ with $card\ I\leq\tau$ in variables $(X_{j})_{j\in J}$ with coefficients in the ultrapower ${\tilde{V}}=\Pi_{\mathcal{U}}V$ has a solution in $\tilde{V}$} if and only if all its finite subsystems have.
Theorem 26.
[…]
Proof.
Let $\mathcal{E}$ be the family of all subfields $k^{\prime}$ of $k$ which are field extensions of finite type over ${\bf F}_{p}$. For such subfield $k^{\prime}$ there exists a lifting $t_{k^{\prime}}:k^{\prime}\rightarrow V$ as we have seen above since $k^{\prime}/{\bf F}_{p}$ is separable generated. We have $k^{\prime}={\bf F}_{p}({\bar{y}}_{k^{\prime}})$ for a finite system of elements ${\bar{y}}_{k^{\prime}}$ from $k$. Let $G_{k^{\prime}}$ be the polynomial equations from ${\bf F}_{p}[Y]$ satisfied by ${\bar{y}}_{k^{\prime}}$ (certainly $G_{k^{\prime}}$ could contain just one equation if we arrange properly ${\bar{y}}_{k^{\prime}}$ but it has no importance for us). The existence of $t_{k^{\prime}}$ says that $G_{k^{\prime}}$ has a solution $y_{k^{\prime}}$ in $V$. If $k^{\prime}_{1}\subset k^{\prime}_{2}$ for some fields $k^{\prime}_{i}\in\mathcal{E}$, $i=1,2$ then the elements ${\bar{y}}_{k^{\prime}_{1}},{\bar{y}}_{k^{\prime}_{2}}$ satisfy some polynomial equations $G_{k^{\prime}_{1}k^{\prime}_{2}}$ from ${\bf F}_{p}[Y]$, several $Y$ which include $G_{k^{\prime}_{i}}$ for $i=1,2$. Certainly the existence of $t_{k^{\prime}_{2}}$ assure us that $G_{k^{\prime}_{1}k^{\prime}_{2}}$ has a solution in $V$. Let $G$ be the union of all possible $G_{k^{\prime}_{1}k^{\prime}_{2}}$ when $k^{\prime}_{1},k^{\prime}_{2}$ runs in $\mathcal{E}$.
Let $U_{1}$ be a set with $card\ U_{1}>card\ k$ and $\mathcal{U}_{1}$ an ultrafilter given by Proposition 25. Then $G$ has a solution in $V_{1}=\prod_{{\mathscr{U}}_{1}}V$ because each finite subsystems of $G$ has already a solution in $V$ given by the lifting of some big enough $k^{\prime}\in\mathcal{E}$. So we get a lifting $t_{k}$ of $k$ in $V_{1}$.
Repeating this procedure we find some sets $(U_{n})$ and some ultrafilters $(\mathcal{U}_{n})$ and define $V_{n}$ and $k_{n}$, so that $k_{n+1}\cong\prod_{{\mathscr{U}}_{n+1}}k_{n}$ and $V_{n+1}=\prod_{{\mathscr{U}}_{n+1}}V_{n}$ with
[…]
The idea is to build ultrafilters ${\mathscr{U}}_{n}$ one by one in such a way that a desired map
$\widetilde{t}\colon\widetilde{k}\rightarrow\widetilde{V}\ \ \mbox{would be the limit of compatible maps }\ \ t_{n}\colon k_{n}\rightarrow V_{n+1}.$
Remark 27.
The proofs of Theorems 23, 26 are similar and in fact we may find $\widetilde{V}$ such that there exist a cross-section $\widetilde{s}:\widetilde{\Gamma}\rightarrow\widetilde{K}^{*}$ and a lifting $\widetilde{t}:\widetilde{k}\rightarrow\widetilde{V}$ in the same time.
Theorem 28.
Let $V$ be a valuation ring containing a field $F$ of positive characteristic. Assume that the residue field $k$ of $V$ is a separable extension of $F$ and one of the following conditions holds
  1. (1)
    $V$ is Henselian,
  2. (2)
    $\dim V=1$ and $V$ is a separable extension of $F$.
Then $V$ is a filtered colimit of complete intersection $F$-algebras.
Proof.
Let $\Gamma$ be the value group of $V$. Let $E$ be a finitely generated $F$-subalgebra (in fact finitely presented by [15, Theorem 4]) of $V$ and $w:E\rightarrow V$ the inclusion. We will show that $w$ factors through a complete intersection $V$-algebra. It is enough to replace $V$ by the intersection of $V$ with the fraction field $K_{E}$ of $E$. Thus we may assume that $\Gamma$ is countable generated and $k$ is countable generated over $F$ by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorems 23, 26 for $V$, or the completion $\hat{V}$ of $V$ in case $\dim V=1$. Thus in both case $V$ is Henselian and so $\widetilde{V}$ is Henselian too. Thus we may assume $\widetilde{k}\subset\widetilde{V}$ (see Theorem 26 and Remark 27) and so it is enough to apply Theorem 24. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so through a complete intersection $F$-algebra $D$ since $\widetilde{k}/k$ and $k/F$ are separable.
Thus $w$ or the composite map $E\rightarrow V\rightarrow\hat{V}$ when $\dim V=1$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$, respectively $\hat{V}$ when $\dim V=1$.
Remains the case 2), that is $\dim V=1$. Since $V$ is separable over $F$ we see that $w(H_{E/F})V\not=0$, where $H_{E/F}$ is the ideal defining the non smooth locus of $E$ over $F$ as in Lemma 11. Then $w$ factors also by $D$ by Lemma 20. The proof ends applying [28, Lemma 1.5].
Corollary 7 has the following version when $\Gamma$ is not necessarily free.
Corollary 29.
Let $V$ be a valuation ring containing its residue field $k$ with characteristic $p>0$. Then $V$ is a filtered colimit of some complete intersection $k$-algebras of finite type. Moreover, $V$ is a filtered colimit of some complete intersection ${\bf F}_{p}$-algebras of finite type.